Find the length of each side of the rectangle so that it has the given perimeter.
width 2x-5
P= 50in

Answer by tutorcecilia(2144)

Use the formula for the perimeter of a rectangle:
P=2(length+width)
Plug-in the values and solve for the missing term (x)
P=50
Width = 2x-5
Length=x
.
P=2(length+width)
50=2[x+(2x-5)]
50=2(x+2x-5)
50=2(3x-5)
50/2=2/2(3x-5)
25=3x-5
25+5=3x
30=3x
30/3=3x/3
10=x (length)
.
Width = 2x-5=2(10)-5=20-5=15
.
check by plugging all of the values back into the original formula:
P=2(length+width)
50=2(10+15)
50=2(25)
50=50 [checks out]